Q Search this course ork Module 1 Homework Ass gam-Score 50 00% Save Subnst Assignment for Grading Check My Work (3 remaining) Click here to read the eBook: Market Value Ratios P/E AND STOCK PRICE Ferrell nc. recenty reported net income of $5 mimion. It has 260,000 shares of common stock, which currently trades at $51 now, its net income will be $7.5 million. Over the next year, it also Assuming Ferrelfs price/earnings ratio remains at its current level, what will be its stock price 1 year from now? Do not round intermediate a share. Ferrell continues to expand and anticipates that I year from anticipates issuing an additional 65,000 shares of stock so that 1 year from now it will have 32,000 shanes of common stock te calculations. Round your answer to the nearest cent. $
Answer #1

For current period,

PE ratio = Market Cap/Net income

PE Ratio = (51*260000)/5000000

PE Ratio = 2.652

Fore next year,

Net Income = 7.5 million

Total Number of shares outstanding = 325,000

PE Ratio = 2.652

PE ratio = Market Cap/Net income

Market Cap = 2.652 * 7500000

Price per share = (2.652*7500000)/325000

Price per share = $61.2